OpenStudy (precal):

Why is the solution for this inequality all real numbers? I don't get it 9 times absolute value (-x+5)+16 is greater than or equal to 16

OpenStudy (anonymous):

\(9|-x+5|+16 \ge 16\) \(9|-x+5| \ge 0\) \(|-x+5| \ge 0\) Absolute value is always positive, so all real numbers have an absolute value greater than or equal to 0

OpenStudy (precal):

Thanks you are awesome.